OpenStudy (anonymous):

1% of 500 is what number

OpenStudy (babalooo):

1% is the same as the fraction 1/100. what is 1/100 of 500?

OpenStudy (anonymous):

50/500 =1% therefore the answer is 50

OpenStudy (anonymous):

\[\frac{ 1 }{ 100 } \times \frac{ x }{ 500 }\]

OpenStudy (anonymous):

cross multiply

OpenStudy (anonymous):

100 * x = 100x 1 * 500 = 500

OpenStudy (anonymous):

divide 100x = 500 ----------- 100 100 x = 5% :)

OpenStudy (babalooo):

i second that and it should be a = sign not a x for the equation on top. not 50 @scbierste

OpenStudy (anonymous):

It should be an equal sign, I was simply showing the 'cross multiply' concept. 10% of 500 would have been 50.

OpenStudy (anonymous):

Either way, hope I was of help.

OpenStudy (anonymous):

@miraclemilktea2 your right, i just looked at the answer that i put up.. it should equal .01 not.1 woops.. my bad

OpenStudy (anonymous):

@miraclemilktea2 are you any good at calc?

OpenStudy (anonymous):

Depends on what the problem is, no promises in my guessing that I can ensure help to you. xD

OpenStudy (anonymous):

Determine the points at which the function f(x)=ln(|7x−32|) is discontinuous

OpenStudy (anonymous):

that shold be Determine the points at which the function f(x)=ln(|7x−32|) is discontinuous

OpenStudy (anonymous):

7x-32 inside the abs. mark

OpenStudy (anonymous):

I'm not sure about that one but heres a website that may be able to help you if you haven't already looked upon it. http://archives.math.utk.edu/visual.calculus/1/continuous.5/ :)

OpenStudy (anonymous):

perfect!!! thanks :]